OpenStudy (anonymous):

Form a polynomial f(x) with real coefficients having the given degrees and zeros. a.) Degree 5; Zeros: -6,-i,-9+1 b.) Degree 4; Zeros: 5; multiplicity: 2; 2i c.)Degree 4; zeros: 5-5i; -1; multiplicity 2

OpenStudy (anonymous):

@satellite73 Could you help me with this problem please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

which one first /

OpenStudy (anonymous):

the first one please

OpenStudy (anonymous):

.) Degree 5; Zeros: -6,-i,-9+1 i assume there is a typo

OpenStudy (anonymous):

degree 5 zeros \(-6,-i,-9+i\) is my guess

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

let me see sorry one sec

OpenStudy (anonymous):

yeah you're right

OpenStudy (anonymous):

ok since one zero is \(-6\) one factor is \(x+6\)

OpenStudy (anonymous):

since one zero is \(-i\) the other one is its conjugate \(i\) and the factors are \[(x+i)(x-i)=x^2+1\]

OpenStudy (anonymous):

the hardest part if finding a quadratic with zeros \(-9+i\) and \(-9-i\) but there are 3 ways to do it one is hard, one is easy and one is really really easy

OpenStudy (anonymous):

lets do the easy way first work backwards starting with \[x=-9+i\] add \(9\) to get \[x+9=i\] square (carefully) and get \[(x+9)^2=i^2\\ x^2+18x+81=-1\] and finally \[x^2+18x+19\]is your quadratic

OpenStudy (anonymous):

lots of typos here let me fix them

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

quadratic is \[x^2+18x+82\]

OpenStudy (anonymous):

on account of \(81+1=82\) not \(19\) so you have to multiply out \[(x+6)(x^2+1)(x^2+18x+82)\] i would cheat

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%28x%2B6%29%28x^2%2B1%29%28x^2%2B18x%2B82%29 you can check that this is right by looking at the zeros

OpenStudy (anonymous):

now we can find the quadratic the real real easy way if you like, when we do C

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

want to do B first?

OpenStudy (anonymous):

it is the easiest of the three

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

Degree 4; Zeros: 5; multiplicity: 2; 2i

OpenStudy (anonymous):

one factor is \(x-5\) but since the "multiplicity" is 2 that means the factor is \((x-5)^2\)

OpenStudy (anonymous):

the other two are \(2i\) and \(-2i\) so the quadratic is \[(x+2i)(x-2i)=x^2+4\] and you only have to multiply out \[(x-5)^2(x^2+4)\]

OpenStudy (anonymous):

c.)Degree 4; zeros: 5-5i; -1; multiplicity 2

OpenStudy (anonymous):

\(-1\) with multiplicity 2 means one factor is \((x+1)^2\)

OpenStudy (anonymous):

now lets find the quadratic with zeros at \(5-5i\) and \(5+5i\) the real easy way it requires memorizing something if the zero is \(a+bi\) then the quadratic is \[x^2-2ax+a^2+b^2\] in your case \(a=5,b=5\) so the quadratic is \[x^2-2\times 5x+5^2+5^2\] or \[x^2-10x+50\]

OpenStudy (anonymous):

final job is \[(x+1)^2(x^2-10x+50)\]

OpenStudy (anonymous):

it says it's wrong?

OpenStudy (anonymous):

wait nvm sorry

OpenStudy (anonymous):

Thank you so much! You really helped me a lot! :) @satellite73

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

do you think you have time to help me with one more problem? Well it's one problem but it has like 3 sub problems to it

OpenStudy (anonymous):

go ahead and post in a new thread, i will take a look

OpenStudy (anonymous):

okay thankyou!